OpenStudy (anonymous):

find all the solutions of: 2sinxcosx+2sinx-cosx-1=0

OpenStudy (anonymous):

2 sinx {cosx +1) = 1 + cosx

OpenStudy (anonymous):

@dgamma3 nw can u solve

OpenStudy (anonymous):

good

OpenStudy (anonymous):

come on.....@dgamma3 just try!!

OpenStudy (anonymous):

trying bro...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

2 sinx {cosx +1) = 1 + cosx

OpenStudy (anonymous):

2 sinx = 1 sin x = 1/2

OpenStudy (anonymous):

nw find the general solution

OpenStudy (anonymous):

if sin x = sin y \[x = n \Pi + (-1)^ny\]

OpenStudy (anonymous):

@dgamma3 did u understand!

OpenStudy (anonymous):

umm, i got the sin x = 1/2. x = inverse sin (1/2)?

OpenStudy (anonymous):

not really sure where the whole sin x = sin y came from though, soz

OpenStudy (anonymous):

u have nt studied this

OpenStudy (anonymous):

no lol, this was a revision question. studied this 2 yrs ago.

OpenStudy (anonymous):

thanks for your help though

OpenStudy (anonymous):

ok..........so x = sin inverse 1/2

OpenStudy (anonymous):

when sin will be 1/2 i mean which angle

OpenStudy (anonymous):

that should be ur answer

OpenStudy (anonymous):

30 degrees?

OpenStudy (anonymous):

yup pie/6

OpenStudy (anonymous):

in radian measure

OpenStudy (anonymous):

much appreciated.

OpenStudy (anonymous):

so does this really mean, for every 1 unit up and sqrt{3} horiztonal, 2 units horizontal?

OpenStudy (anonymous):

i mean, 2 units diagonal

OpenStudy (anonymous):

assuming cos is 2/sqrt(3)